The product of two numbers is 20. If one of the numbers is 8/13 what is the other number?
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

The product of two numbers is 20. If one of the numbers is 8/13 what is the other number?
-----
(8/13)x = 20
------
x = (13/8)20
x = 260/8
x = 32.5
------------
Cheers,
Stan H.
-------------

Answer by Alan3354(69443)

The product of two numbers is 20. If one of the numbers is 8/13 what is the other number?
-----------
a*b = 20
a = 8/13
----
(8/13)*b = 20
Solve for b